OpenStudy (anonymous):

A bag contains 3 red marbles, 4 blue marbles, and 5 white marbles. What is the probability of pulling out a white marble, and then a blue marble if none are replaced?

OpenStudy (nubeer):

P (WB) = 5/12 * (4/11)

OpenStudy (anonymous):

What is the probability of pulling out a white marble??

OpenStudy (anonymous):

@chuck83 answer me!

OpenStudy (anonymous):

i'm the one asking for help!

OpenStudy (anonymous):

i asked a part of it

OpenStudy (anonymous):

okay do u know how to calculate the probability ?@chuck83

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

Probability = (number of events favourable)/(number of total events) for example in a die, probability of getting even number if, we want to find then Counting the favourable events what are the favourable events = getting 2,4,6 so there are 3 favourable outcomes total events are 1,2,3,4,5,6 , so there are 6 events on total So the probability of getting even number is 3/6 that gives us 0.5 [ remember that probability never exceeds 1 as in the extreme cases Favourable events can be zero as in " getting 7 on the 6 sided dice" or favourable events can be total events as in " getting any number on 6 sided dice"]

OpenStudy (anonymous):

Thats for how to calculate Probability basically.

OpenStudy (anonymous):

In your problem given "A bag contains 3 red marbles, 4 blue marbles, and 5 white marbles. What is the probability of pulling out a white marble, and then a blue marble if none are replaced?" If none are replaced the number of marbles available for first picking would be 3R+4B+5W = 3+4+5 = 12 marbles in the first picking one marble is taken out so the availble marbles are 12-1 = 11 now probability of picking white marble: favourable events are 5 as there are 5 marbles of white color and total marblesare 12 so the probability of pulling out a white marble = 5/12 now probability of pulling out a blue marble: favourable events are 4 as there are 4 marbles of blue color total marbles now become 12-1 = 11 so probability of pulling out a blue marble = 4/11 so probability on the whole is (5/12)*(4/11) = 20/132

OpenStudy (anonymous):

@chuck83 got it?